# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

### v1

@@ Abstract.
Stochastic policies are widely used in continuous-time reinforcement learning algorithms. However, executing a stochastic policy and evaluating its performance in a continuous-time environment remain open challenges. This work introduces and rigorously analyzes a policy execution framework that samples actions from a stochastic policy at discrete time points and implements them as piecewise constant controls. We prove that as the sampling mesh size tends to zero, the controlled state process converges weakly to the dynamics with coefficients aggregated according to the stochastic policy. We explicitly quantify the convergence rate based on the regularity of the coefficients and establish an optimal first-order convergence rate for sufficiently regular coefficients. Additionally, we show that the same convergence rates hold with high probability concerning the sampling noise, and further establish a $1/2$-order almost sure convergence when the volatility is not controlled. Building on these results, we analyze the bias and variance of various policy evaluation and policy gradient estimators based on discrete-time observations. Our results provide theoretical justification for the exploratory stochastic control framework in [H. Wang, T. Zariphopoulou, and X.Y. Zhou, J. Mach. Learn. Res., 21 (2020), pp. 1-34].
@@ 1. Introduction
Reinforcement learning (RL) has catalyzed significant changes across various domains by attacking complex sequential decision-making challenges in a data-driven way. The theoretical framework for analyzing RL problems and designing RL algorithms has been predominantly based on the discrete-time Markov decision processes [5, 30]. However, many practical applications require real-time interaction with environments where time progresses continuously, such as autonomous driving, stock trading, and vehicle routing. Recent studies [33, 38, 20, 12] have shown that certain algorithms originally designed for discrete-time RL can become unstable when applied in high-frequency settings. This highlights the need for a new theoretical framework for designing RL algorithms that are robust to time discretization.
@@
Exploratory control and stochastic policy
Recently, [36] proposed a novel exploratory control formulation with stochastic policies for continuous-time RL problems, which has served as the foundation for designing various RL algorithms; see, e.g., [18, 19, 20, 15, 8, 11, 37, 12, 29, 32].
This control formulation in [36] has three key features: (1) The agent incorporates stochastic policies into the control problem, which models exploration in RL. These policies are mappings from the state space to distributions over the action space, aligning with the notion of “relaxed control” in classical control theory. (2) The agent continuously samples actions from a stochastic policy, and the state process evolves according to a new diffusion process, with its coefficients integrated with respect to the sampling distribution specified by the stochastic policy. This auxiliary state dynamic is often referred to as the exploratory or aggregated dynamics. (3) The optimal policy remains stochastic due to the inclusion of an additional entropy regularizer in the cost objective, in contrast to the deterministic/strict policies typically found in standard control problems.
@@
However, despite its success in algorithm design, the use of stochastic policies in continuous-time RL problems introduces theoretical and practical gaps. Specifically, the connection between exploratory dynamics with aggregated coefficients and state dynamics controlled by frequent sampling from a stochastic policy remains unjustified. In fact, interacting with a diffusion process by continuously sampling from a stochastic policy requires generating uncountably many independent random variables, which raises measure-theoretical issues and renders the resulting controlled state dynamics ill-defined (see the end of Section 2 for more details).
To address this issue, [32] proposes approximating exploratory dynamics by sampling actions from a stochastic policy only at discrete time points. This approach naturally aligns with the practical implementation of stochastic policies. However, for algorithms designed based on aggregated dynamics, executing a stochastic policy in discrete time introduces a bias that depends on the sampling frequency. This leads to the following natural question:
What is the precise impact of sampling frequency on the performance of a stochastic policy and the corresponding learning algorithms?
To the best of our knowledge, this question has only been addressed in the context of linear-quadratic (LQ) RL problems with Gaussian policies [12, 32]. In this special setting, due to the LQ structure, the analysis is simplified by examining the discrepancy between the first and second moments of the aggregated dynamics and the controlled dynamics with sampled actions. Unfortunately, these techniques developed specifically for LQ RL problems are clearly not applicable to general RL problems. In particular, it is essential to quantify the difference between the entire distributions of two state processes.
@@
This paper analyzes the accuracy of the discrete-time execution of stochastic policies for general diffusion processes and quantifies its impact on the associated learning algorithms.
  • We show that when actions are sampled from a stochastic policy at a discrete time grid $\mathscr{G}$ and executed as a piecewise constant control process, the resulting state process, referred to as the sampled dynamics, converges weakly to the aggregated dynamics as the time stepsize approaches zero. In particular, under sufficient regular coefficients, we show that the convergence rate is of first order $O(|\mathscr{G}|)$, with $|\mathscr{G}|$ being the time stepsize (Theorem 4.1). This matches the optimal convergence rate for standard piecewise constant control approximations [16]. We further quantify the precise weak convergence rate for cases with less regular coefficients (Theorem 4.2).
@@
  • We prove that the sampled dynamics, in fact, converges almost surely with respect to the randomness in policy execution. Specifically, we show that the weak convergence order to the aggregated dynamics holds with high probability with respect to the randomness used to sample actions (Theorem 4.3). We further show that when the volatility is uncontrolled, the sampled dynamics converges almost surely to the aggregated dynamics with an order of $\mathcal{O}(|\mathscr{G}|^{1/2})$, for each fixed realization of the Brownian motion (Theorem 4.4). The analysis leverages concentration inequalities for bounded and sub-exponential martingale difference sequences to control the concentration behavior of the sampling errors in terms of the sampling frequency (see the discussions below Theorems 4.3 and 4.4).
    The almost sure convergence result indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time, i.e., the convergence can be achieved by refining the sampling frequency while fixing the realization of the system noise. Note that such an almost sure convergence fails when the volatility is controlled, as shown in Example 4.6.
@@
Our approaches and most related works
This paper quantifies the accuracy of the discrete-time execution of a stochastic policy and its impact on continuous-time RL algorithms. The accuracy of policy execution has been studied in [14, 32] for LQ problems with Gaussian policies. However, as mentioned earlier, the techniques developed in [14, 32] are not applicable to RL problems with general state dynamics and stochastic policies considered in this paper.
For general state dynamics, the weak convergence of sampled dynamics has recently been studied by [4, 3]. They adopt a different definition of sampled dynamics, treating it as being driven by a random measure, and establish the weak convergence of the state process without quantifying the convergence order. In contrast to the time-varying control considered in [4, 3], we analyze sampled dynamics controlled by a piecewise constant control process. Moreover, we take a completely different approach and establish the weak convergence rate under various regularity conditions by leveraging the solution regularity of the associated partial differential equations (PDEs). Beyond weak convergence, we also establish convergence in the almost sure sense with respect to the policy execution noise. These strong convergence results require a careful analysis of the concentration properties of the state dynamics concerning the policy execution noise.
In addition to discrete sampling error, we also analyze the error incurred when approximating continuous-time RL algorithms using discrete-time observations. [18] analyzes a similar error but assumes that observations are generated from the aggregated dynamics. [2] analyzes both errors for LQ problems with uncontrolled volatility.
@@
Our analysis of the policy execution error is fundamentally different from classical numerical analysis of time discretizations for stochastic differential equations (SDEs). In particular, numerical analysis of SDEs focuses on the error caused by discretizing the system noise, such as Brownian motion, while both the SDE and its time-discretized version have the same coefficients. In our setting, we treat the controlled SDE as a black-box environment evolving continuously and focus on the error introduced by the piecewise constant execution of a stochastic policy. As a result, the environment noise is not discretized, but the limiting SDE (the aggregated dynamics) and the sampled dynamics have different coefficients. The former involves aggregated coefficients, while the latter has random coefficients determined by the specific procedure used to sample the policy.
@@
The rest of the paper is organized as follows. Section 2 briefly reviews the use of stochastic policies in continuous-time RL. Section 3 presents the problem setup and introduces the sampled dynamics rigorously. Section 4 presents the main convergence results of the sampled dynamics, and Section 5 applies them to analyze continuous-time RL algorithms. Finally, we conclude in Section 6. The proofs of all statements are included in Section 7.
@@ 2. Preliminary
In this section, we review the exploratory control formulation and the use of stochastic policies for continuous-time RL algorithms. The precise assumptions and setup can be found in Section 3.
@@ 2.1. Classical stochastic control
Recall that in a classical stochastic control problem, an agent aims to control the state process and optimize an objective functional through an adapted control process taking values in a Polish space $A$. Specifically, given an admissible control process $\{a_{t}\}_{t\geq 0}$, the corresponding state process $\{X_{t}\}_{t\geq 0}$ is governed by the following dynamics: [^1] To simplify the presentation, we assume that the state processes have a deterministic initial condition. However, our analysis and results extend naturally to cases with a random initial condition, provided it is independent of the Brownian motion $W$ and satisfies appropriate integrability conditions. We also assume that the Brownian motion’s dimension equals the state variable’s dimension; otherwise, we can introduce more Brownian motions or state variables to make them equal.
@@
where $x_{0}\in{\mathbb{R}}^{d}$ is a given initial state, $b:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d\times d}$ are given drift and volatility coefficients, respectively, and $W$ is an $d$-dimensional Brownian motion supported on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. The agent’s objective is to maximize the following objective
@@
over all admissible controls $\{a_{t}\}_{t\geq 0}$, where $r:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$ is a given instantaneous reward function, and $h:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a given terminal reward function.
It is known that under suitable conditions (see [39]), an optimal control for (2.2) exists and is given by a feedback form, i.e., there exists a measurable function $\bm{u}:[0,T]\times\mathbb{R}^{d}\to A$ such that the optimal control is obtained by $a_{t}=\bm{u}(t,X_{t}^{\bm{u}})$ for all $t\in[0,T]$, where $\{X_{t}^{\bm{u}}\}_{t\geq 0}$ satisfies the following dynamics:
@@
The function can be determined using the coefficients $b,\sigma,r,$ and $h$. Once determined, the optimal control can be implemented by evaluating ${\bm{u}}$ at the current state of the system.
2.2. RL, stochastic policy and exploratory control
@@
To avoid getting trapped in local optima, a key feature of RL is explicitly exploring the environment through randomized actions. A common approach is to use stochastic policies $\bm{\pi}:[0,T]\times{\mathbb{R}}^{d}\to\mathcal{P}(A)$, which maps the current state to a probability distribution over the action space. Given a stochastic policy $\bm{\pi}$, the agent interacts with the environment by sampling noisy actions from $\bm{\pi}$ and observing the resulting system states. Over time, the agent refines the stochastic policy to optimize the objective.
To describe the performance of a stochastic policy $\bm{\pi}$, [36] suggests that the associated state process $\tilde{X}$ evolves according to the following dynamics (called the exploratory dynamics therein):
@@
$\mathbb{E}^{W}\left[\int_{0}^{T}\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t+h(\widetilde{X}_{T})\right],\quad\textnormal{with $\tilde{r}(t,x)\coloneqq\int_{A}r(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. }$ (2.6)
The dynamics (2.4) and the objective (2.6) are derived heuristically in [36] under the assumption that the agent interacts with the system (2.1) by continuously sampling independent random actions from the policy $\bm{\pi}$. As a result of these random actions, a “law-of-large-numbers” heuristic suggests that the infinitesimal change in the state at time $t$ is determined by the coefficients in (2.5), aggregated over the sampling distribution defined by $\bm{\pi}$. Likewise, the instantaneous reward is also aggregated with respect to this sampling distribution. In the sequel, we refer to (2.4) as the aggregated dynamics to highlight this aggregating effect.
However, the rigorous justification of the connection between (2.4) and the state dynamics resulting from the adoption of a stochastic policy $\bm{\pi}$ in (2.1), as well as the relationship between (2.6) and the original reward functional (2.2), remains to be established. Indeed, interacting with the SDE (2.1) by continuously sampling from a stochastic policy creates measure-theoretical issues. As already pointed out in [32, Remark 2.1], it is impossible to construct a family of non-constant random variables $(\xi_{t})_{t\in[0,1]}$ such that $(\xi_{t})_{t\in[0,1]}$ is (essentially) pairwise independent and $t\mapsto\xi_{t}$ is Lebesgue measurable. This implies that if one controls (2.1) by continuously generating independent actions, the resulting coefficients are not progressively measurable, rendering the conventional stochastic integral ill-defined; see [3] for more discussion. Moreover, it is practically infeasible to generate these independent samples continuously.
Therefore, both from theoretical and practical perspectives, evaluating the performance of a stochastic policy $\bm{\pi}$ requires discretely sampling actions from the policy and applying them to (2.1), as suggested by [32]. This work quantifies the accuracy of such a discrete execution of a stochastic policy and its impact on the corresponding RL algorithms.

### v2

@@ Abstract.
Stochastic policies (also known as relaxed controls) are widely used in continuous-time reinforcement learning algorithms. However, executing a stochastic policy and evaluating its performance in a continuous-time environment remain open challenges. This work introduces and rigorously analyzes a policy execution framework that samples actions from a stochastic policy at discrete time points and implements them as piecewise constant controls. We prove that as the sampling mesh size tends to zero, the controlled state process converges weakly to the dynamics with coefficients aggregated according to the stochastic policy. We explicitly quantify the convergence rate based on the regularity of the coefficients and establish an optimal first-order convergence rate for sufficiently regular coefficients. Additionally, we prove a $1/2$-order weak convergence rate that holds uniformly over the sampling noise with high probability, and establish a $1/2$-order pathwise convergence for each realization of the system noise in the absence of volatility control. Building on these results, we analyze the bias and variance of various policy evaluation and policy gradient estimators based on discrete-time observations. Our results provide theoretical justification for the exploratory stochastic control framework in [H. Wang, T. Zariphopoulou, and X.Y. Zhou, J. Mach. Learn. Res., 21 (2020), pp. 1-34].
@@ 1. Introduction
Reinforcement learning (RL) has catalyzed significant changes across various domains by attacking complex sequential decision-making challenges in a data-driven way. The theoretical framework for analyzing RL problems and designing RL algorithms has been predominantly based on the discrete-time Markov decision processes [35]. However, many practical applications require real-time interaction with environments where time progresses continuously, such as autonomous driving, stock trading, and vehicle routing. Recent studies [38, 44, 22, 13] have shown that certain algorithms originally designed for discrete-time RL can become unstable when applied in high-frequency settings. This highlights the need for a new theoretical framework for designing RL algorithms that are robust to time discretization.
@@
Stochastic policy in continuous-time RL
Recently, [42] introduced a novel exploratory control framework for continuous-time RL that fundamentally relies on the use of stochastic policies. Stochastic policies are mappings from the state space to distributions over the action space, aligning with the notion of “relaxed control” in classical control theory. By sampling actions from these stochastic policies, the agent assigns positive probabilities to different actions at each state, unlike the deterministic/strict policies typically employed in standard control problems. This approach effectively captures the concept of exploration in RL, which is essential for the agents to learn about the environment. [^1] Exploration in RL broadly refers to deliberately acting suboptimally to gather information about the environment. Stochastic policies achieve this by randomizing actions, but exploration can also be driven by methods such as adding exploration bonuses to encourage actions in less-certain outcomes; see [35, Chapter 2]. Since then, the use of stochastic policies and the corresponding exploratory control framework have become a foundational basis for developing a wide range of RL algorithms for continuous-time problems; see, e.g., [20, 21, 22, 16, 9, 12, 43, 47, 13, 37, 34].
@@
However, despite its popularity in algorithm design, the use of stochastic policies in continuous-time RL presents both theoretical and practical challenges. In particular, the behavior of the state dynamics under frequent sampling from a stochastic policy has not been rigorously analyzed.
In [42], the authors assume that agents continuously sample from a stochastic policy and heuristically argue that the resulting dynamics follow a new diffusion process, with coefficients aggregated according to the sampling distribution of the policy. This aggregated dynamics offers a mathematically convenient framework for designing continuous-time RL algorithms. However, it requires generating uncountably many independent random variables, which is computationally infeasible and raises measure-theoretical concerns, rendering the controlled dynamics ill-defined. As a consequence, such an aggregated dynamics are not directly implementable. (see footnote 3 for details).
In this paper, we consider a natural and practical approach to implement stochastic policies. Rather than sampling actions continuously, we sample them at discrete time points and apply them to the continuous-time state system. As a result, the state dynamics are governed by a piecewise constant control process, whose evolution differs from the aggregated dynamics studied in [42].
For algorithms developed under the aggregated dynamics, such discrete-time execution of stochastic policies introduces a bias that must be carefully quantified.
In summary, this paper aims to address the following two questions:
Q1:
Does the aggregated dynamics in [42] capture the state dynamics induced by frequently sampling a stochastic policy? If so, in what sense, and what is the resulting error?
Q2:
What is the precise impact of the sampling frequency of a stochastic policy on the performance of the corresponding learning algorithms?
To the best of our knowledge, these questions have only been studied in the context of linear-quadratic (LQ) RL problems with Gaussian policies [13, 37]. In this special setting, due to the LQ structure, the analysis is simplified by examining the discrepancy between the first and second moments of the aggregated dynamics and the controlled dynamics with sampled actions. Unfortunately, these techniques developed specifically for LQ RL problems are clearly not applicable to general RL problems. In particular, it is essential to quantify the difference between the entire distributions of two state processes.
@@
This paper analyzes the convergence of the discrete-time execution of stochastic policies for general diffusion processes and quantifies its impact on the associated learning algorithms.
  • We show that when actions are sampled from a stochastic policy at a discrete time grid $\mathscr{G}$ and executed as a piecewise constant control process, the resulting state process, referred to as the sampled dynamics, converges weakly to the aggregated dynamics as the time stepsize approaches zero. In particular, under sufficiently regular coefficients, we show that the convergence rate is of first order $O(|\mathscr{G}|)$, with $|\mathscr{G}|$ being the time stepsize (Theorem 4.1). This matches the optimal convergence rate for standard piecewise-constant control approximations with smooth coefficients [18]. We further quantify the precise weak convergence rate for cases with less regular coefficients (Theorem 4.2).
@@
  • We prove that the sampled dynamics, in fact, converges almost surely with respect to the randomness in policy execution. Specifically, we show that the sampled dynamics converge at order $\mathcal{O}(|\mathscr{G}|^{1/2})$, averaged over the system noise but uniformly across the realizations of the sampling noise with high probability (Theorem 4.3). We further show that when the volatility is uncontrolled, the sampled dynamics converge at order $\mathcal{O}(|\mathscr{G}|^{1/2})$ for almost every realization of the system noise (Theorem 4.4). We show in Example 4.7 that the condition of uncontrolled volatility is necessary for the convergence of the sampled dynamics for each fixed system noise.
    These almost sure convergence results allow for interpreting the convergence of the sampled dynamics to the aggregated dynamics as a law of large numbers in time, i.e., convergence is achieved by increasing the sampling frequency without averaging over multiple realizations of the sampling noises at given time points. See Corollary 4.1 and Remark 4.3 for its implications on the improved sample complexity of associated Monte Carlo estimators.
@@
Our approaches and key technical challenges
Our convergence rate analysis for the discrete-time execution of stochastic policies requires addressing several technical issues beyond those encountered in existing literature. These challenges arise because the sampled dynamics contain two distinct sources of randomness: one from control randomization and the other from the underlying system noise.
@@
Specifically, the sampled dynamics have random coefficients due to control randomization, and they converge to a limiting diffusion process with a different set of deterministic coefficients, resulting from aggregation over the stochastic policy. Quantifying this convergence rate has not been addressed in the literature, and it is fundamentally different from the classical analysis of time discretizations for stochastic differential equations (SDEs) (see, e.g., [25]). Specifically, standard numerical analysis of SDEs focuses on the error caused by discretizing the system noise, while both the SDE and its time-discretized version share the same coefficients. In our setting, the environment noise is not discretized; instead, the error arises from the control randomization.
Moreover, we quantify the convergence rate of the sampled dynamics under different criteria, including the weak error with respect to both system and sampling noises, the weak error conditioned on the sampling noise, and almost sure convergence with respect to both noises. Identifying precise conditions to ensure convergence under these criteria requires carefully distinguishing the roles of system and sampling noises and employing different analytical techniques, including the regularity theory of partial differential equations (PDEs) and various martingale concentration inequalities.
In particular, when quantifying the weak error conditioned on the sampling noise (Theorem 4.3), we leverage the averaging over the system noise and apply the Azuma–Hoeffding inequality for bounded martingales. For the almost sure convergence with respect to both noises (Theorem 4.4), more precise estimates involving conditional Orlicz norms of sub-exponential martingales are required to carefully characterize the concentration behavior with respect to both system and sampling noises. To the best of our knowledge, this is the first work that applies concentration inequalities to the convergence rate analysis of sampling schemes for SDEs.
Most related works
This paper quantifies the accuracy of the discrete-time execution of a stochastic policy and the performance of continuous-time RL algorithms based on discrete-time observations. The accuracy of policy execution has been studied in [15, 37] for LQ problems with Gaussian policies; however, their techniques do not extend to the RL problems with general state dynamics and stochastic policies as considered here. Regarding the performance of continuous-time RL algorithms using discrete-time observations, most existing works assume that the observations are generated directly from the desired continuous-time state dynamics, independent of time discretizations (see, e.g., [2, 20, 48, 49]). This assumption no longer holds in the present setting with stochastic policies, since the agent observes the sampled dynamics, which differ from their continuous-time limit.
For general state dynamics and general stochastic policies, the weak convergence of sampled dynamics has recently been studied in [4, 3]. In their setting, the policy is evaluated continuously at the current state, and weak convergence is established by interpreting the sampled action as a random measure, though without providing a convergence rate. In contrast, we consider a more practical setting where the policy is evaluated only at discrete time points, so that the sampled dynamics are governed by a piecewise constant control process. Moreover, we establish not only the weak convergence rate, leveraging the regularity theory of PDEs, but also the high-probability bound and almost sure convergence rate with respect to the policy execution noise.
For state dynamics with drift controls only, [7] establishes the strong convergence of time-discretized sampled dynamics under the assumption that the drift coefficient is bounded. This boundedness condition is restrictive, as it rules out the linear dynamics commonly encountered in many RL problems [2, 15, 37]. Our result removes this limitation by allowing drift coefficients with linear growth. To address the challenges posed by unbounded coefficients, we employ advanced sub-exponential concentration inequalities with conditional Orlicz norms to quantify the convergence rate.
Organization of the paper
@@
Section 2 briefly reviews the use of stochastic policies in continuous-time RL. Section 3 presents the problem setup and introduces the sampled dynamics rigorously. Section 4 presents the main convergence results of the sampled dynamics, and Section 5 applies them to analyze continuous-time RL algorithms. Section 6 presents the proofs of the main results. Appendix A contains the proofs of useful Lemmas and intermediary results.
@@ 2. Preliminary
In this section, we review the use of stochastic policies in continuous-time RL problems, while avoiding unnecessary technical details. The precise assumptions and setup are given in Section 3.
@@ 2.1. Classical stochastic control
Recall that in a classical stochastic control problem, an agent aims to control the state process and optimize an objective functional through a non-anticipative control process taking values in a Polish space $A$. Specifically, given a non-anticipative control process $\{a_{t}\}_{t\geq 0}$, the corresponding state process $\{X_{t}\}_{t\geq 0}$ is governed by the following dynamics: [^2] To simplify the presentation, we assume that the state processes have a deterministic initial condition. The analysis and results extend naturally to cases with a random initial condition that is independent of the Brownian motion $W$ and satisfies appropriate integrability conditions. We also assume without loss of generality that the Brownian motion and the state variable have the same dimensions.
@@
where $x_{0}\in{\mathbb{R}}^{d}$ is a given initial state, $b:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d\times d}$ are given drift and volatility coefficients, respectively, and $W$ is a $d$-dimensional Brownian motion supported on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. The agent’s objective is to maximize the following objective
@@
over all non-anticipative controls, where $r:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$ is a given instantaneous reward function, and $h:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a given terminal reward function. We refer the reader to [45, Chapter 2, Section 4] for a more detailed formulation of stochastic control problems.
It is known that under mild regularity conditions (see e.g., [32, 26]), an optimal control of (2.2) exists and can be expressed in feedback form. That is, there exists a function $\bm{u}:[0,T]\times\mathbb{R}^{d}\to A$, called a policy, such that the optimal control is obtained by $a_{t}=\bm{u}(t,X_{t}^{\bm{u}})$ for all $t\in[0,T]$, where $\{X_{t}^{\bm{u}}\}_{t\geq 0}$ satisfies the following dynamics:
@@
The policy $\bm{u}$} can be determined using the coefficients $b,\sigma,r,$ and $h$. Once determined, the optimal control can be implemented by evaluating $\bm{u}$} at the current state of the system.
2.2. Stochastic policy in RL and the associated state dynamics
@@
To avoid getting trapped in local optima, a key feature of RL is explicitly exploring the environment. A common approach is to use stochastic policies $\bm{\pi}:[0,T]\times{\mathbb{R}}^{d}\to\mathcal{P}(A)$, which maps the current state to a probability distribution over the action space. Given a stochastic policy $\bm{\pi}$, the agent interacts with the state by sampling noisy actions from $\bm{\pi}$ and observing the resulting state transitions. Over time, the agent refines the stochastic policy to optimize the objective.
A crucial step in designing learning algorithms with stochastic policies is to characterize the performance of a stochastic policy $\bm{\pi}$. [42] suggests that given the policy $\bm{\pi}$, the associated state process $\widetilde{X}$ evolves according to the following dynamics:
@@
$\mathbb{E}\left[\int_{0}^{T}\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t+h(\widetilde{X}_{T})\right],\quad\textnormal{with $\tilde{r}(t,x)\coloneqq\int_{A}r(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. }$ (2.6)
In the sequel, we call (2.4) the aggregated dynamics to emphasize that its coefficients are obtained by averaging the coefficients of (2.1) with respect to the sampling distribution defined by $\bm{\pi}$.
The dynamics (2.4) and the objective (2.6) are derived heuristically in [42] under the assumption that the agent interacts with the system (2.1) by continuously sampling independent random actions from the policy $\bm{\pi}$. Under this assumption, the authors determine the state coefficients (2.5) and the instantaneous reward (2.6) by examining the first and second moments of the infinitesimal state change and applying a “law-of-large-numbers” heuristic. However, the aggregated dynamics (2.4) differs from the practically observed state process, since realizing it would require continuously sampling independent actions. [^3] In fact, interacting with the SDE (2.1) by continuously sampling from a stochastic policy creates measure-theoretical issues. As pointed out in [37, Remark 2.1], it is impossible to construct a family of non-constant random variables $(\xi_{t})_{t\in[0,1]}$ such that $(\xi_{t})_{t\in[0,1]}$ is (essentially) pairwise independent and $t\mapsto\xi_{t}$ is Lebesgue measurable. This implies that if one controls (2.1) by continuously generating independent actions, the resulting coefficients are not progressively measurable, rendering the conventional stochastic integral ill-defined.
A more natural and practical implementation of a stochastic policy is to sample actions only at discrete time points and apply them to the state process (2.1) (see [13, 37]). The resulting state dynamics is governed by (2.1) with a piecewise constant control process, whose evolution differs from the aggregated dynamics (2.4). In the sequel, we show that these state dynamics converge to the aggregated dynamics (2.4) as the sampling frequency increases, and we quantify the corresponding convergence rates. We also analyze the impact of this approximation on algorithms designed based on (2.4).